### v1

@@ 6 Physical properties: model & photometry dependence
The SED fitting technique introduces several systematic uncertainties in derived properties coming from the assumptions made about the models, which may lead to a disagreement in the estimated stellar masses of up to a factor of (e.g Maraston et al. 2006; Kannappan & Gawiser 2007; Conroy 2013; Lower et al. 2020; Pacifici et al. 2023). In this Section, we consider how our choices of the SED modules (see Table 3) influence the main physical properties (stellar masses and SFRs). We validate the impact of the number of assumptions made about the IMF, fixed metallicity, SSP models, SFH prescription, dust attenuation and emission laws, and AGN models (see Table 9 for a list of the changed parameters given in App. D). To quantify the effect, we make only a change of one parameter with respect to the default configuration (see Table 3). To discriminate the impact of the incorporation of the WISE photometry, we compare the physical properties estimated with or without WISE photometry or only with W1 and W2 bands in Sec. 6.4. We finally compare the CIGALE derived SFR with the ones estimated from the emission line measurement in Sec. 6.5. To preserve the computational time, we validate the influence of the model assumptions on a smaller sample of galaxies representing seven main galaxy classes: 1,750 BL AGN, 3,900 NL AGN, 8,393 composite, 8,819 star-forming, 8,526 passive, 8,846 retired and 9,962 other galaxies. The selection of this representative sample is described in the App. D.1.
@@ 6.1 Stellar Mass: Choice of the stellar components
We consider three main IMF choices: i) Salpeter (1955) assuming a power-law distribution with a heavier slope towards high-mass stars, ii) Kroupa (2001) assuming a broken power-law IMF with a shallower slope at higher stellar masses, implying fewer massive stars compared to Salpeter (1955), iii) Chabrier (2003), which combines a log-normal distribution for low-mass stars and a power-law distribution for higher masses and thus is considered as a reasonable middle ground between Salpeter (1955) and Kroupa (2001). The Chabrier (2003) IMF typically results in lower stellar mass estimates compared to the Salpeter (1955) IMF, because the Chabrier (2003) IMF reduces the contribution of high-mass stars, which are more massive and contribute more to the total stellar mass. To mitigate the difference in stellar mass estimates we commonly assume a conversion factor between Chabrier (2003) and other IMFs, namely:
@@
as shown by e.g. Cimatti et al. (2008); Longhetti & Saracco (2009); Bolzonella et al. (2010); Ilbert et al. (2010). Here, we take the opportunity, to assess whether a simple scaling is sufficient and whether it varies for different types of galaxies. We find a constant median difference of -0.24 dex (see Sec. 5.1 for the definition of the median difference and Table 5 for the quantitative comparison of the estimates with different model assumptions) between stellar masses obtained assuming Chabrier (2003) and Salpeter (1955) stellar masses. This median difference is in agreement with the commonly used conversion factor and is independent of the stellar mass and galaxy class.
@@
To generate the DESI VAC we rely on the Bruzual & Charlot (2003) models. We discriminate the degree to which our estimates are different from the ones derived under the assumption of the Maraston (2005) model. The main difference between Maraston (2005) and Bruzual & Charlot (2003) relies on the incorporation of the thermally pulsating asymptotic giant branch (TP-AGB) stars in the evolution of galaxies in Maraston (2005) models. TP-AGB stars are evolved stars that can significantly affect the integrated light of galaxies, particularly in the NIR wavelength range. Due to the inclusion of TP-AGB stars, the Maraston (2005) model might yield higher stellar mass estimates compared to Bruzual & Charlot (2003) for galaxies with significant contributions from these stars, especially at intermediate and old ages (see more details in e.g Maraston 2005; Conroy & Gunn 2010; Kriek et al. 2010; Maraston & Strömbäck 2011). While both models (BC03 and M05) include remnants, the exact mass contributions from remnants might not differ drastically between the two models. However, any differences would stem from the specific stellar evolution prescriptions and the IMF adopted in each model. The way these effects are included are not the same even for a simple Salpeter 1955 IMF (Maraston, 2005) and some offsets just from that cannot be excluded (e.g. Maraston et al., 2013). Typically, remnants constitute a relatively small fraction of the total stellar mass, so while there might be differences, they should not be as pronounced as the differences arising from the TP-AGB treatment. We note that recent work using James Webb Telescope spectra reports the spectroscopic detection of the TP-AGB in galaxies at high-z and models with little contribution from that phase do not to fit the data well (Lu et al., 2024). Different works reported an offset of dex between stellar masses of Bruzual & Charlot (2003) and Maraston 2005 (e.g. Ilbert et al., 2010; Pozzetti et al., 2010; Kawinwanichakij et al., 2020). The change of the SSP model to Maraston (2005) introduces a median difference of -0.26 dex in our stellar mass estimates with respect to the ones derived with Bruzual & Charlot (2003) models (see Sec. 5.1 for the definition of the median difference and Table 5 for the quantitative comparison of the estimates with different model assumptions).
@@
Commonly, physical properties are derived relying on the parametric SFHs, although they may suffer from strong biases in recovering the proper SFHs due to the assumption of the simplistic prescription (e.g. Ciesla et al. 2015; Carnall et al. 2019; Lower et al. 2020; Leja et al. 2022). As a response to the necessity of more advanced SFH prescription, non-parametric SFHs were proposed (e.g. Leja et al. 2019; Lower et al. 2020; Ciesla et al. 2023b). We verify how our stellar mass estimates change when using a non-parametric SFH module (sfhNlevels; see Table 9) implemented in CIGALE (Ciesla et al., 2023b). The formula of the non-parametric SFH is based on time bins in which the SFRs are constant and linked together by the bursty continuity (Tacchella et al., 2022) instead of the assumption of the analytical function. We refer the reader to Arango-Toro et al. (2023); Ciesla et al. (2023a, b) for more details about the non-parametric SFH module. For DESI galaxies, the SED fitting of the optical-MIR photometry with non-parametric SFH has a negligible impact on the stellar mass estimates across the entire stellar mass range (with a median difference of 0.03). The median difference is two times higher for AGN and star-forming galaxies () than for passive and retired galaxies (0.02; see Tab. 5).
@@
Stellar metallicity is poorly constrained from photometric data alone due to the age-metallicity-dust degeneracy (e.g. Worthey, 1994; Papovich et al., 2001). To overcome this problem, most commonly SED fitting-based approaches rely on fixing the metallictiy to a solar value (e.g. Małek et al., 2018; Boquien et al., 2019) or leaving it as a free but constant value over the lifetime of the galaxy (e.g. Carnall et al., 2018; Johnson et al., 2021). These assumptions may affect other parameters of interest such as the stellar mass or SFRs introducing mass-dependent systematics (e.g. Pforr et al., 2012; Mitchell et al., 2013; Thorne et al., 2022a). On the other hand, other works (e.g. Osborne & Salim 2024) do not report the dependence of the stellar mass estimates on the choice of metallicity. We also validate the impact on the stellar mass estimates with models for which the metallicity is allowed to vary (see Table 9) instead of assuming the metallicity fixed to a solar value. For our catalog, allowing metallicity to vary introduces a median difference of -0.04 across the entire stellar mass range. The median difference is two times higher for NL AGN and composite () than for the remaining galaxy classes (; see Tab. 5).
@@
| IMF (Chabrier 2003 vs Salpeter 1955) | -0.244 | 0.362 | -0.248 | 0.367 | -0.246 | 0.364 |
| SSP (Bruzual & Charlot 2003 vs Maraston 2005) | -0.258 | 0.382 | -0.235 | 0.348 | -0.262 | 0.388 |
| SFH (delayed with extra burst vs non parametric) | 0.027 | 0.066 | 0.039 | 0.094 | 0.044 | 0.084 |
| Z (metallicity fixed to solar vs variable) | -0.039 | 0.076 | -0.040 | 0.073 | -0.046 | 0.101 |
| DustAtt (Calzetti et al. 2000 vs Charlot & Fall 2000) | -0.031 | 0.063 | -0.024 | 0.063 | -0.109 | 0.167 |
| DustEm (Draine et al. 2014 vs Dale et al. 2014) | -0.000 | 0.018 | -0.000 | 0.028 | -0.000 | 0.049 |
| AGN (Fritz et al. 2006 vs no AGN) | 0.004 | 0.018 | 0.007 | 0.024 | -0.001 | 0.045 |
| AGN (Fritz et al. 2006 vs Stalevski et al. 2012, 2016) | 0.001 | 0.006 | 0.003 | 0.007 | 0.005 | 0.030 |
| SEDs ( vs ) | -0.002 | 0.055 | -0.009 | 0.071 | -0.013 | 0.187 |
| SEDs ( vs ) | 0.036 | 0.166 | 0.067 | 0.190 | 0.044 | 0.212 |
@@ 6.2 Stellar Mass: Choice of the dust models
We consider an alternate dust attenuation model proposed by Charlot & Fall 2000 (see Table 9) to test the systematics. In contrast to Calzetti et al. (2000), this recipe assumes different attenuation for young ( Myr) and old stars ( Myr). Young stars are attenuated in the birth clouds, while both young and old stars are attenuated in the interstellar medium. In CIGALE, the implementation of the Charlot & Fall (2000) law is more flexible, giving the freedom to choose the values of input parameters (attenuation of the ISM, slopes of power-law attenuation curves for the birth cloud, and the ISM, and the ratio of the total attenuation). Both attenuation laws are modeled by a power law and normalized to the attenuation in the V band. The main difference in the shape of attenuation curves appears at , where Charlot & Fall (2000) is flatter than the one given by Calzetti et al. (2000). For example, Mitchell et al. (2013) estimated that the stellar mass can be underestimated by up to 0.6 dex by assuming the Calzetti et al. (2000) for massive galaxies. For DESI galaxies we find a median difference of dex across the entire stellar mass range, but the median difference is higher for AGN and composite galaxies () than for the remaining classes (; see Tab. 5).
We also consider the prescription of the dust emission model given by Dale et al. 2014 (see Table 9), which is much simpler than the complex model of Draine et al. (2014). The star-forming component is described by where is the dust mass heated by the radiation field at intensity and represents the relative contributions of the different local SEDs (Dale et al., 2014). The parameter is the only free parameter and is tightly connected with the color. Due to a limited variation of the PAH with respect to , the model does have problems with proper modeling of the dust in metal-poor galaxies (e.g. Engelbracht et al. 2005). However, when using only optical-MIR SEDs, there is no difference in the used dust emission model as indicated by the zero median difference in the stellar mass estimates using Dale et al. (2014) and Draine et al. (2014) prescriptions (see Sec. 5.1 for the definition of the median difference and Table 5 for the quantitative comparison of the estimates with different model assumptions).
@@ 6.3 Choice of the AGN models
Finally, we consider the impact of incorporating AGN templates on the stellar mass estimates by comparing the estimates from our VAC with estimates from CIGALE when the AGN contribution is fixed to 0 for the representative sample. As suggested by Thorne et al. (2022b), the inclusion of the AGN component is argued to be crucial not to overestimate the light coming from the host galaxies. However, our results suggest that the stellar masses for the general sample of galaxies are not affected significantly by the contribution of AGN as the median difference is low ( = 0.004 for the entire representative sample). On the other hand, the median difference is high for BL AGN (; see Fig. 10). This implies that the incorporation of the AGN models affects the stellar mass estimates (the median difference is slightly higher than the median stellar mass error; ) for the BL AGN, while for remaining galaxy classes this effect is negligible from the statistical point of view (the scatter as shown in Fig. 10 is large if we consider the individual sources). We note that the difference in stellar mass for BL AGN depends on redshift and stellar mass (see Fig. 11), while for the remaining classes, the stellar mass differences are independent of redshift and considered stellar mass range.
The change of the AGN model to the one proposed by Stalevski et al. 2012, 2016 (for the description of AGN models see Sec. 3.4) does not affect the stellar mass estimates for the entire galaxy population (see Table 5) and the difference is the highest for the BL AGN (with a median difference of 0.05 and NMAD of 0.14). However, we note that the scatter for individual galaxies is large; for individual sources, the masses might be higher even for dex when using Fritz et al. (2006) models than when relying on Stalevski et al. (2012, 2016).
@@ 6.4 Stellar Mass: Choice of the photometry
Aside from the model assumptions discussed in the previous Section, the SED coverage is one of the most important ingredients for deciding on the reliability of the SED-derived physical properties. The coverage of the is highly desired to obtain reliable estimates of the contribution from young and old stars and AGN (e.g. Thorne et al. 2022b, a). However, numerous works suggest that robust and reliable stellar mass estimates require only optical photometry to find the tight relation between optical color and stellar M/L ratio (e.g. Bell & de Jong 2001; Bell et al. 2003; Gallazzi & Bell 2009; Zibetti et al. 2009; Taylor et al. 2010). Gallazzi & Bell (2009) showed that stellar mass estimates obtained on one optical color are not biased against the estimates based on the optical-NIR SEDs or spectral features. However, Gallazzi & Bell 2009 based their estimates on the BC03 models that are NIR featureless. Other works instead find that only optical + NIR photometry can break the age/metallicity/dust degeneracy (Maraston et al., 2010).
@@
In this Section, we compare the estimates of stellar mass for DESI galaxies assessing further in the wavelength, i.e. using optical colors alone (grz) with the ones obtained based on the optical-MIR SED fit adding i) only WISE1 and WISE2 () and all four WISE photometry (). We find a negligible impact of incorporating to the SED fit on the stellar mass estimates (; see Table 5) for the entire galaxy population. The discrepancy between stellar mass estimates increases for AGN ( = 0.13 and -0.06 for BL, and NL AGN, respectively, and NMAD 0.26 and 0.16 for BL, and NL AGN, respectively). Independently of the galaxy class, the scatter is large dex, thus the stellar masses for individual sources may differ significantly. The stellar masses obtained only based on grz photometry are still similar to the ones derived based on the grzW14 SED fits ( = 0.04) but the scatter increases (NMAD = 0.17; see Table 5). The degree of difference in stellar mass estimates depends on the galaxy types, and we find that this is higher for the BL AGN (with = 0.20 and NMAD = 0.34) than for the remaining classes (see Fig. 12). The difference in stellar mass is independent of redshift except for the low redshift (), where the stellar mass estimated with WISE photometry are higher than the ones estimated only based on the optical bands (see Fig. 13). Independently on redshift, the stellar masses can differ up to dex for AGN and composite galaxies, and even up to dex for star-forming and passive galaxies.
6.5 SFR: Choice of the models & photometry

### v2

@@ 6 Physical properties: model and photometry dependence
The SED fitting technique introduces several systematic uncertainties in derived properties coming from the assumptions made about the models, which may lead to a disagreement in the estimated stellar masses of up to a factor of (e.g., Maraston et al. 2006; Kannappan & Gawiser 2007; Conroy 2013; Lower et al. 2020; Pacifici et al. 2023). In this section, we consider how our choices of the SED modules (see Table 3) influence the main physical properties (stellar masses and SFRs). We validate the impact of the number of assumptions made about the IMF, fixed metallicity, SSP models, SFH prescription, dust attenuation and emission laws, and AGN models (see Table 9 for a list of the changed parameters given in Appendix D). To quantify the effect, we made only one change, namely, in a parameter related to the default configuration (see Table 3). To discriminate the impact of the incorporation of the WISE photometry, we compared the physical properties estimated with or without WISE photometry or only with W1 and W2 bands in Sect. 6.4. Finally, we compared the CIGALE derived SFR with the ones estimated from the emission line measurement in Sect. 6.5. To preserve the computational time, we validated the influence of the model assumptions on a smaller sample of galaxies representing seven main galaxy classes: 1,750 BL AGNs, 3,900 NL AGNs, 8,393 composite, 8,819 star-forming, 8,526 passive, 8,846 retired, and 9,962 other galaxies. The selection of this representative sample is described in Appendix D.1.
@@ 6.1 Stellar mass: choice of the stellar components
We considered three main IMF choices: i) Salpeter (1955) assuming a power-law distribution with a heavier slope towards high-mass stars, ii) Kroupa (2001) assuming a broken power-law IMF with a shallower slope at higher stellar masses, implying fewer massive stars compared to Salpeter (1955), iii) Chabrier (2003), which combines a log-normal distribution for low-mass stars and a power-law distribution for higher masses and thus is considered as a reasonable middle ground between Salpeter (1955) and Kroupa (2001). The Chabrier (2003) IMF typically results in lower stellar mass estimates compared to the Salpeter (1955) IMF, because the Chabrier (2003) IMF reduces the contribution of high-mass stars, which are more massive and have a greater contribution to the total stellar mass. To mitigate the difference in stellar mass estimates we commonly assume a conversion factor between Chabrier (2003) and other IMFs, namely:
@@
as shown by, for instance, Cimatti et al. (2008), Longhetti & Saracco (2009), Bolzonella et al. (2010), and Ilbert et al. (2010). Here, we take the opportunity, to assess whether a simple scaling is sufficient and whether it varies for different types of galaxies. We find a constant median difference of -0.24 dex (see Sect. 5.1 for the definition of the median difference and Table 5 for the quantitative comparison of the estimates with different model assumptions) between stellar masses obtained assuming Chabrier (2003) and Salpeter (1955) stellar masses. This median difference is in agreement with the commonly used conversion factor and is independent of the stellar mass and galaxy class.
@@
To generate the DESI VAC, we relied on the Bruzual & Charlot (2003) models. We discriminate the degree to which our estimates are different from the ones derived under the assumption of the Maraston (2005) model. The main difference between Maraston (2005) and Bruzual & Charlot (2003) relies on the incorporation of the thermally pulsating asymptotic giant branch (TP-AGB) stars in the evolution of galaxies in Maraston (2005) models. TP-AGB stars are evolved stars that can significantly affect the integrated light of galaxies, particularly in the NIR wavelength range. Due to the inclusion of TP-AGB stars, the Maraston (2005) model might yield higher stellar mass estimates compared to Bruzual & Charlot (2003) for galaxies with significant contributions from these stars, especially at intermediate and old ages (see more details in e.g., Maraston 2005; Conroy & Gunn 2010; Kriek et al. 2010; Maraston & Strömbäck 2011). While both models (BC03 and M05) include remnants, the exact mass contributions from remnants might not differ drastically between the two models. However, any differences would stem from the specific stellar evolution prescriptions and the IMF adopted in each model. The way these effects are included are not the same even for a simple Salpeter 1955 IMF (Maraston, 2005) and some offsets just from that cannot be excluded (e.g., Maraston et al., 2013). Typically, remnants constitute a relatively small fraction of the total stellar mass, so while there might be differences, they should not be as pronounced as the differences arising from the TP-AGB treatment. We note that recent works using James Webb Telescope spectra reports the spectroscopic detection of the TP-AGB in galaxies at high-z and models with little contribution from that phase do not fit the data well (Lu et al., 2024). Different works have reported an offset of dex between stellar masses of Bruzual & Charlot (2003) and Maraston 2005 (e.g., Ilbert et al., 2010; Pozzetti et al., 2010; Kawinwanichakij et al., 2020). The change of the SSP model to Maraston (2005) introduces a median difference of -0.26 dex in our stellar mass estimates with respect to the ones derived with Bruzual & Charlot (2003) models. We refer to Sect. 5.1 for the definition of the median difference and Table 5 for the quantitative comparison of the estimates with different model assumptions.
@@
Typically, physical properties are derived relying on the parametric SFHs, although they may suffer from strong biases in recovering the proper SFHs due to the assumption of the simplistic prescription (e.g., Ciesla et al. 2015; Carnall et al. 2019; Lower et al. 2020; Leja et al. 2022). As a response to the necessity of more advanced SFH prescription, non-parametric SFHs have been proposed (e.g., Leja et al. 2019; Lower et al. 2020; Ciesla et al. 2023). We verified how our stellar mass estimates change when using a non-parametric SFH module (sfhNlevels; see Table 9) implemented in CIGALE (Ciesla et al., 2023). The formula of the non-parametric SFH is based on time bins in which the SFRs are constant and linked together by the bursty continuity (Tacchella et al., 2022), instead of the assumption of the analytical function. We refer the reader to Arango-Toro et al. (2023); Ciesla et al. (2024, 2023) for more details about the non-parametric SFH module. For DESI galaxies, the SED fitting of the optical-MIR photometry with non-parametric SFH has a negligible impact on the stellar mass estimates across the entire stellar mass range (with a median difference of 0.03). The median difference is two times higher for AGNs and star-forming galaxies () than for passive and retired galaxies (0.02; see Table 5).
@@
Stellar metallicity is poorly constrained from photometric data alone due to the age-metallicity-dust degeneracy (e.g., Worthey, 1994; Papovich et al., 2001). To overcome this problem, most commonly SED fitting-based approaches rely on fixing the metallictiy to a solar value (e.g., Małek et al., 2018; Boquien et al., 2019) or leaving it as a free but constant value over the lifetime of the galaxy (e.g., Carnall et al., 2018; Johnson et al., 2021). These assumptions may affect other parameters of interest, such as the stellar mass or SFRs introducing mass-dependent systematics (e.g., Pforr et al., 2012; Mitchell et al., 2013; Thorne et al., 2022a). On the other hand, other works (e.g., Osborne & Salim 2024) do not report the dependence of the stellar mass estimates on the choice of metallicity. We also validated the impact on the stellar mass estimates with models where the metallicity is allowed to vary (see Table 9), instead of assuming the metallicity fixed to a solar value. For our catalog, allowing metallicity to vary introduces a median difference of -0.04 across the entire stellar mass range. The median difference is two times higher for NL AGNs and composite () than for the remaining galaxy classes (; see Table 5).
@@
| IMF (Chabrier 2003 vs. Salpeter 1955) | -0.244 | 0.362 | -0.248 | 0.367 | -0.246 | 0.364 |
| SSP (Bruzual & Charlot 2003 vs. Maraston 2005) | -0.258 | 0.382 | -0.235 | 0.348 | -0.262 | 0.388 |
| SFH (delayed with extra burst vs. non parametric) | 0.027 | 0.066 | 0.039 | 0.094 | 0.044 | 0.084 |
| Z (metallicity fixed to solar vs. variable) | -0.039 | 0.076 | -0.040 | 0.073 | -0.046 | 0.101 |
| DustAtt (Calzetti et al. 2000 vs. Charlot & Fall 2000) | -0.031 | 0.063 | -0.024 | 0.063 | -0.109 | 0.167 |
| DustEm (Draine et al. 2014 vs. Dale et al. 2014) | -0.000 | 0.018 | -0.000 | 0.028 | -0.000 | 0.049 |
| AGN (Fritz et al. 2006 vs. no AGN) | 0.004 | 0.018 | 0.007 | 0.024 | -0.001 | 0.045 |
| AGN (Fritz et al. 2006 vs. Stalevski et al. 2012, 2016) | 0.001 | 0.006 | 0.003 | 0.007 | 0.005 | 0.030 |
| SEDs ( vs. ) | -0.002 | 0.055 | -0.009 | 0.071 | -0.013 | 0.187 |
| SEDs ( vs. ) | 0.036 | 0.166 | 0.067 | 0.190 | 0.044 | 0.212 |
@@ 6.2 Stellar mass: Choice of the dust models
We considered an alternate dust attenuation model proposed by Charlot & Fall 2000 (see Table 9) to test the systematics. In contrast to Calzetti et al. (2000), this recipe assumes different attenuation for young ( Myr) and old stars ( Myr). Young stars are attenuated in the birth clouds, while both young and old stars are attenuated in the interstellar medium. In CIGALE, the implementation of the Charlot & Fall (2000) law is more flexible, giving the freedom to choose the values of input parameters (attenuation of the ISM, slopes of power-law attenuation curves for the birth cloud, and the ISM, and the ratio of the total attenuation). Both attenuation laws are modeled by a power law and normalized to the attenuation in the V band. The main difference in the shape of attenuation curves appears at , where Charlot & Fall (2000) is flatter than the one given by Calzetti et al. (2000). For example, Mitchell et al. (2013) estimated that the stellar mass can be underestimated by up to 0.6 dex by assuming the Calzetti et al. (2000) for massive galaxies. For DESI galaxies we find a median difference of dex across the entire stellar mass range, but the median difference is higher for AGN and composite galaxies () than for the remaining classes (; see Table 5).
We also considered the prescription of the dust emission model given by Dale et al. 2014 (see Table 9), which is much simpler than the complex model of Draine et al. (2014). The star-forming component is described by where is the dust mass heated by the radiation field at intensity and represents the relative contributions of the different local SEDs (Dale et al., 2014). The parameter is the only free parameter and is tightly connected with the color. Due to a limited variation of the PAH with respect to , the model does have problems with proper modeling of the dust in metal-poor galaxies (e.g., Engelbracht et al. 2005). However, when using only optical-MIR SEDs, there is no difference in the used dust emission model, as indicated by the zero median difference in the stellar mass estimates using Dale et al. (2014) and Draine et al. (2014) prescriptions (see Sect. 5.1 for the definition of the median difference and Table 5 for the quantitative comparison of the estimates with different model assumptions).
@@ 6.3 Choice of the AGN models
Finally, we considered the impact of incorporating AGN templates on the stellar mass estimates by comparing the estimates from our VAC with estimates from CIGALE when the AGN contribution is fixed to 0 for the representative sample. As suggested by Thorne et al. (2022b), the inclusion of the AGN component is argued to be crucial not to overestimate the light coming from the host galaxies. However, our results suggest that the stellar masses for the general sample of galaxies are not affected significantly by the contribution of AGN as the median difference is low ( = 0.004 for the entire representative sample). On the other hand, the median difference is high for BL AGN (; see Fig. 10). This implies that the incorporation of the AGN models affects the stellar mass estimates (the median difference is slightly higher than the median stellar mass error; ) for the BL AGNs; whereas for the remaining galaxy classes, this effect is negligible from the statistical point of view (i.e., the scatter shown in Fig. 10 is large if we consider the individual sources). We note that the difference in stellar mass for BL AGNs depends on redshift and stellar mass (see Fig. 11), while for the remaining classes, the stellar mass differences are independent of redshift and considered stellar mass range.
The change of the AGN model to the one proposed by Stalevski et al. 2012, 2016 (for the description of AGN models; see Sect. 3.4) does not affect the stellar mass estimates for the entire galaxy population (see Table 5) and the difference is the highest for the BL AGNs (with a median difference of 0.05 and NMAD of 0.14). However, we note that the scatter for individual galaxies is large; for individual sources, the masses might be higher even for dex when using Fritz et al. (2006) models than when relying on Stalevski et al. (2012, 2016).
@@ 6.4 Stellar mass: choice of the photometry
Aside from the model assumptions discussed in the previous section, the SED coverage is one of the most important ingredients for deciding on the reliability of the SED-derived physical properties. The coverage of the is highly desired to obtain reliable estimates of the contribution from young and old stars and AGNs (e.g., Thorne et al. 2022b, a). However, numerous works suggest that robust and reliable stellar mass estimates require only optical photometry to find the tight relation between optical color and stellar M/L ratio (e.g., Bell & de Jong 2001; Bell et al. 2003; Gallazzi & Bell 2009; Zibetti et al. 2009; Taylor et al. 2010). Gallazzi & Bell (2009) showed that stellar mass estimates obtained on one optical color are not biased against the estimates based on the optical-NIR SEDs or spectral features. However, Gallazzi & Bell 2009 based their estimates on the BC03 models that are NIR featureless. Other works instead find that only optical + NIR photometry can break the age, metallicity, and dust degeneracy (Maraston et al., 2010).
@@
In this section, we compare the estimates of stellar mass for DESI galaxies assessing further in the wavelength, namely, using optical colors alone (grz) with the ones obtained based on the optical-MIR SED fit, adding i) only WISE1 and WISE2 () and all four WISE photometry (). We find a negligible impact when incorporating to the SED fit on the stellar mass estimates (; see Table 5) for the entire galaxy population. The discrepancy between stellar mass estimates increases for AGNs ( = 0.13 and -0.06 for BL and NL AGNs, respectively, and NMAD 0.26 and 0.16 for BL, and NL AGNs, respectively). Independently of the galaxy class, the scatter is large dex, thus the stellar masses for individual sources may differ significantly. The stellar masses obtained only based on grz photometry are still similar to the ones derived based on the grzW14 SED fits ( = 0.04), but the scatter increases (NMAD = 0.17; see Table 5). The degree of difference in stellar mass estimates depends on the galaxy types, and we find that this is higher for the BL AGNs (with = 0.20 and NMAD = 0.34) than for the remaining classes (see Fig. 12). The difference in the stellar mass is independent of redshift except for the low redshift (), where the stellar mass estimated with WISE photometry are higher than the ones estimated only based on the optical bands (see Fig. 13). Independently of redshift, the stellar masses can differ up to dex for AGNs and composite galaxies – and even up to dex for star-forming and passive galaxies.
6.5 SFR: Choice of the models and photometry